Black Sea Construction Company has a contract to construct a $6,000,000 oil rig at an estimated cost of $5,300,000. The contract is to start in July 2017, and the oil rig is to be completed in October 2019. The following data pertain to the construction period.

2017    2018    2019
Costs to date $1,325,000 $3,720,000 $6,300,000
Estimated costs to complete 3,975,000 2,480,000
Progress billings during the year 1,200,000 3,200,000 1,600,000
Cash collected during the year 1,000,000 2,340,000 2,660,000


What amount of gross profit (loss) should Black Sea recognize in 2018 using the percentage-of-completion method?
Answer #1

Answer

  • Under Percentage of Completion method,
    For 2018, there will be GROSS (LOSS) of $ (900,000)
  • Working for above answer

Working

2017

2018

2019

A

Contract Price

$               5,300,000.00

$               5,300,000.00

$                    5,300,000.00

B

Cost Incurred to Date

$                1,325,000.00

$               3,720,000.00

$                    6,300,000.00

C

Estimated cost yet to be incurred to complete the contract

$                3,975,000.00

$               2,480,000.00

$                                         -  

D = B+C

Total Cost

$                5,300,000.00

$               6,200,000.00

$                    6,300,000.00

E = (B/D) x 100

% of Completion

25.00%

60.00%

100.00%

F = A x E

Revenue to date

$                1,325,000.00

$               3,180,000.00

$                    5,300,000.00

G

Revenue of Previous year

$                                     -  

$               1,325,000.00

$                    3,180,000.00

H = F - G

Net Revenue this year

$                1,325,000.00

$               1,855,000.00

$                    2,120,000.00

I (=B)

Cost to date

$                1,325,000.00

$               3,720,000.00

$                    6,300,000.00

J

Cost to date of previous year

$                                     -  

$               1,325,000.00

$                    3,720,000.00

K = I - J

Net Cost for the year

$                1,325,000.00

$               2,395,000.00

$                    2,580,000.00

L = A - D

Estimated Gross Profit (Loss)

$                                     -  

$                 (900,000.00)

$                  (1,000,000.00)

M

Gross Profit (Loss) to be recognised

$                                     -  

$                 (900,000.00) ANSWER

$                     (100,000.00)